at the moment john is 5 years older than mary. in seven years mary age will be 3/4 of john age how old are they now?
Answer by lwsshak3(11628)

at the moment john is 5 years older than mary. in seven years mary age will be 3/4 of john age how old are they now?
**
let x=mary's age now
x+5=john's age now
7 years hence:
x+7=mary's age
x+5+7=x+12=john's age
..
x+7=(3/4)(x+12)
4x+28=3x+36
x=36-28=8
x+5=13
mary's age now=8
john's age now=13